A researcher is interested in estimating the average monthly salary of bank managers in NCR. He wants to be 90% confident that his estimate is correct. If the standard deviation is 8,500 pesos, how large a sample is needed te get the desired information and to be accurate within 1,150 pesos?


1
Expert's answer
2022-04-12T16:29:50-0400

"E=z_{0.05}\frac{\sigma}{\sqrt{n}}"

"n=(\frac{z_{0.05}\sigma}{E})^2=(\frac{1.645*8500}{1150})^2=148."